# Assembly Theory Reduced to Shannon entropy and Rendered Redundant by Naive Statistical Algorithms

### v7

@@ 4 Methods
For the assembly index (Ai), we will explore the proposed counterexamples as presented in [1] with the purpose of countering trivially fallacious arguments that:
@@
    in order for two compression schemes to belong to the $\mathbf{LZ}$ family, they must necessarily be identical;
@@
    in the general case (by accounting for arbitrary strings beyond a handful of cherry-picked examples), the expected compression rate (or expected number of factors in their respective parsings) would not be correlated at the asymptotic limit (as the length of strings/objects increases), even though they may not actually be for an a priori chosen set of particular examples (especially for short sequences).
@@
For the Molecular Assembly Index (MA) introduced in [5] as an application of Ai within the context of molecular complexity, we show that it is not only highly correlated to the LZW of InChI strings, but also highly correlated to the length of the InChI strings themselves. Besides, we show that one of the main benefits of the Ai (‘molecular assembly’ or MA, being experimentally measurable using mass spectrometry data) is also observed for the LZW of InChI strings. See also Section 4.3.
@@
In order to trivially check that (a) is false, one simply needs to note that it is always possible to pick particular cases of strings for which the parsing according to LZ77 differs from the one according to LZ78 [15]. For example, one can see in [1, Figure $3$] that the authors omit the dictionary built for the right square (i.e., the one for the parsing according to the assembly index calculation method). Specifically, this assembly coding dictionary is
[…]
$\left(2,\left[1,1\right]\right),$
$\left(3,\left[2,2\right]\right),$
$\left(4,\left[3,3\right]\right)$
@@
and the final object would be given by the codeword list $\left[4,4\right]$ as indicated by the right square in [1, Figure $3$]. Notice that the set of vertices of the minimum rooted assembly subspace of the final object is given by the set $\left\{1,2,3,4,\left[4,4\right]\right\}$ composed of pointers; and the assembly index is the size of this dictionary excluding the basis element under pointer $1$, that is, Ai $=5-1=4$. Another fact overlooked by the authors is that if one uses an $\mathbf{LZ}$ factorisation/parsing scheme straightforwardly defined by adding to the dictionary the longest prefix of the subsequent subsequence that has already occurred, or a single symbol (in our case, “z”) if there is no such non-empty prefix, then one would obtain exactly the desired parsing
[…]
@@
that the authors of AT overlooked, a parsing which gives the exact assembling process in the minimum rooted assembly subspace of the final object in Equation (7). Notice that the number of factors/codewords in Equation (7) (minus $1$) is the assembly index of the final object. Except for the initial basis element “z”, the parsing in Equation (7) exactly corresponds to the sequence of edge labels of the rooted assembly path of the final object. In this trivial example, we see that the shortest rooted assembly path and the longest one (which belong to the minimum rooted assembly subspace) are the same.
@@
The second part (b) of the counterargument constitutes an intersection of the “Texas sharp shooter” with the “hasty generalisation” fallacies: where one takes the global properties as always the same as the properties resulting from analyses of particular local cases, which they would invariably be, given that the criteria for the latter analyses were chosen to produce the desired local properties.
@@
From the equivalence between Ai and dictionary-based compression schemes formally demonstrated in [4], we know that the set of vertices/objects of the minimum subspace encodes/compresses the final object (in our case, the input string). This holds whether the basis elements are included or not, since the compression rate obtained in the former case or the latter case differs only by an independent constant. Thus, the number of codewords in the list (which results from the parsing of the string according to the assembly index method) is equivalent to the length of the longest (rooted) assembling paths in the minimum subspace, except for those that are basis elements.
@@
For example, for the input string “abracadabra”, a dictionary obtained for the Assembly Theory calculations would be:
[…]
@@
[…]
@@
[…]
while the condeword list is
@@
In the case of LZW/LZ78 we know that the size in bits of the compressed form of the input string (in the asymptotic limit) grows under $c_{1}(n)\mathbf{O}\left(\log\left(c_{1}(n)\right)\right)$ [15, 35], where $c_{1}(n)$ is the size of the LZW/LZ78 dictionary for a string of length $n$. The same also holds for the compressed form (with size $K_{Ai}\left(x\right)$ in bits) of the input string $x$ according to the assembly index calculation method, so that
$K_{Ai}\left(x\right)\leq c_{2}(n)\mathbf{O}\left(\log\left(c_{3}(n)\right)\right)\leq c_{2}(n)\mathbf{O}\left(\log\left(c_{2}(n)\right)\right)\text{ ,}$ (12)
where $c_{2}(n)$ is the length of the list of codewords that constitutes the string $x$ and $c_{3}(n)=Ai(n)$ is the assembly index which is the size of the Ai dictionary [4] for a string $x$ of length $n=\left|x\right|$ with $Ai(n)\leq c_{2}(n)$.
Sequences with larger values of $c_{i}\left(\cdot\right)$, i.e., the number of distinct phrases or codewords in the dictionary, to which the pointers necessarily recur less often, would have lower probability. This is the same key idea used to demonstrate the optimality of $\mathbf{LZ}$ schemes toward the limit given by the noiseless coding theorem; and it is the same idea behind the assembly index as a complexity measure, which the authors claim is a novelty.
In this paper, we further investigate the relationships between these values with experiments that generalise the purported counterexamples to much longer sequences of different configurations, as explained in Section 4.2, thus demonstrating that (b) is indeed such a fallacy.
4.1 The Block Decomposition Method
The Block Decomposition Method (BDM), as introduced in 2013 [36, 37, 7], is an estimator of algorithmic information redundancies $\mathbf{K}\left(\cdot\right)$ defined by a decomposition of the object into parts for which one already has an algorithmic complexity estimation, obtained by means of, for example, the Coding Theorem Method (CTM) [8] based on Algorithmic Probability [38, 39] (AP) and the related Universal Distribution [40] which takes into consideration all the statistical and algorithmic regularities and redundancies in data. The Coding Theorem Method or CTM consists in finding all the formal-theoretic explanations for an object shorter than the object itself by exhaustive means [7]. Given that CTM can therefore only find small programs, BDM extends the power of CTM by stitching these programs together to offer a causal explanation of the object as a test of (non-)randomness while one can always find tighter lower bounds on AP or upper bounds on $\mathbf{K}$ by running CTM further [7]. AP is the probability that a random explanation (e.g. a computer program) can explain an object [40] under a (computable) generative model that ultimately can, by computational universality, be proven to converge, thanks to the Invariance Theorem [15, 38, 41, 39], independent of computable model (up to a ‘small’ constant that has been proven to be stable and small [42, 43]).
The BDM of an object $x$, according to the partition $i$, that is, one of the ways to decompose the object $x$ to which one assigns the corresponding index $i$, is defined by
$BDM(x,i,m)=\sum\limits_{\left(r_{j},n_{j}\right)\in P_{i}\left(x\right)}\big{(}\log(n_{j})+K_{m}(r_{j})\big{)}\text{ ,}$ (13)
@@ 4.1 The Block Decomposition Method
  • $P_{i}\left(x\right)$ is the set of pairs $\left(r_{j},n_{j}\right)$ obtained when decomposing the object $x$ according to a partition $i$ of contiguous parts $r_{j}$, each of which appears $n_{j}$ times in such a partition (i.e., $n_{j}$ is the multiplicity of $r_{j}$), that is, the number of exact repetitions.
@@
Notice that Equation (13) can be expressed for many particular cases, e.g. in the case of unidimensional objects (such as bitstrings) but also any other object, including non-binary strings and n-dimensional objects such as graphs, matrices, images, vectors and tensors (any encodable, computer representable data) where $m$ indicates the method given by the pre-computed table of complexity estimations derived from CTM [7, 34, 42, 32].
Also notice that if one takes a partition $i_{0}$ for which the object’s size is the partition size (or, equivalently, there is only one block/part which is the object itself), then $BDM(x,i_{0},m)$ becomes equivalent to the approximation method indexed as $m$, e.g. the CTM. That is, there is $m$ such that
@@
where $Q$ is a computable (proper) subset of $\left\{\mathbf{M}\colon\exists w^{\prime}\left(\mathbf{M}\left(\emptyset\right)=w^{\prime}\right)\right\}$ and $Q(x)\subset\left\{\mathbf{M}\colon\mathbf{M}\left(\emptyset\right)=x\right\}$ a computable subset of $Q$ that the method $m$ calculates, where the algorithmic probability (AP) of $x$ is approximated by
@@
and $\mathbf{M}\left(x\right)=y$ denotes algorithm $\mathbf{M}$ returning $y$ with input $x$. Deriving from this equivalence, one of the properties of BDM is that it is a complexity measure that approaches $\mathbf{K}\left(\cdot\right)$ in the best case scenario, and approaches block (Shannon) Entropy in the worst case [7], that is,
$\mathbf{K}\left(x\right)\leq\mathbf{O}(1)+BDM(x,i_{0},m)\leq\mathbf{O}(1)+BDM(x,i,m^{\prime})\leq\mathbf{O}(1)+N_{i}\,\mathbf{H}_{i}\left(X^{i}\right)\text{.}$ (15)
The left inequalities of Equation (15) always hold, from the basic properties of algorithmic complexity. From classical information theory, the rightmost inequality holds for sufficiently large $\left|x\right|$ and any computable compression method $m^{\prime}$ that is capable of matching (or surpassing) the statistical compression/encoding achieved by, e.g., Huffman or LZW for each codeword, where
$\mathbf{H}_{i}\left(X^{i}\right)=-\sum\limits_{\left(r_{j},n_{j}\right)\in P_{i}\left(x\right)}p\left(r_{j}\right)\log\left(p\left(r_{j}\right)\right)$ (16)
is the block (Shannon) Entropy $\mathbf{H}_{i}$ of an i.i.d. source $X^{i}$ such that $p\left(r_{j}\right)\to\frac{n_{j}}{N_{i}}$ as $\left|x\right|\to\infty$, the random variable $X^{i}$ can assume values in the set $\left\{r_{1},\dots,r_{j},\dots,r_{\left|P_{i}\left(x\right)\right|}\right\}$, and $N_{i}=\left(\sum\limits_{\left(r_{j},n_{j}\right)\in P_{i}\left(x\right)}n_{j}\right)$.
Thus, in case $m^{\prime}$ is a sufficiently powerful computable approximation and $i$ a partition closer to the partition $i_{0}$ (i.e., the one for which the object’s size is the partition size), one obtains $\mathbf{K}\left(x\right)\sim BDM(x,i,m^{\prime})$; otherwise, in case the object $x$ is indeed generated by a stochastic source (e.g., following an i.i.d. probability distribution of the blocks/parts $r_{j}$), one obtains $BDM(x,i,m^{\prime})\leq\mathbf{O}(1)+N_{i}\,\mathbf{H}_{i}\left(X^{i}\right)$ approaching the equality the more statistically random (i.e., for more uniform probability distributions with $n_{j}\to 1$) the source is. Also note that, as demonstrated in [7] for regular bidimensional partitions, an algorithm can always losslessly retrieve $x$ from the information in the parts/blocks calculated in $BDM(x,i,m^{\prime})$ by an encoding scheme $_{BDM}$ such that
|$x\right>_{BDM}\right|\leq\mathbf{O}(1)+BDM(x,i,m^{\prime})+\mathbf{O}\left(N_{i}\log\left(\left|P_{i}(x)\right|\right)\right)$ (17)
@@
This means that for objects generated according to a probability distribution sufficiently close to the uniform distribution and a sufficiently powerful $m^{\prime}$, BDM can always achieve a statistical lossless compression of the same order as $\mathbf{H}$, while the value of $BDM(x,i,m^{\prime})$ can be smaller than $N_{i}\,\mathbf{H}_{i}\left(X^{i}\right)$. In general, for any object and any type of source (stochastic or otherwise), one can take BDM for partition $i_{0}$ and the only differential factor to achieve a better compression than $\mathbf{H}$ will be the compression method $m$, which can always be the one indexing e.g. Huffman’s, or better approximations to $\mathbf{K}$. Therefore, $\mathbf{H}$ is a (statistical) compression method subsumed within BDM (see also Section 2.6).
A false argument repeated by the authors of Assembly Theory is that approximations to algorithmic Kolmogorov complexity are uncomputable, unlike assembly theory [23]. Not only is this false, as BDM is fully computable [44, 45], even if it can always improve upper and lower bounds on AP and $\mathbf{K}$ by running CTM further [8] to improve on any $\mathbf{H}$ estimation that may falsely indicate high randomness where causal content exists [7]; but also, the assembly index itself is an approximation to $\mathbf{K}$ equivalent to $\mathbf{H}$ and $\mathbf{LZ}$ compression, as we have proven [4], therefore a loose upper bound on $\mathbf{K}$ itself. Hence, the assembly index offers no advantage.
BDM can also deal with stochasticity, by exploring all possible computer program pathways that assemble an object, as fully explained in the Supplementary Information appended to this paper.

### v10

@@ 4 Methods
We investigate the counterexamples provided in [4] with the purpose of countering what we have identified as fallacious arguments and, misleading and incomplete experiments, including:
@@
    in order for two compression schemes to belong to the LZ family, they must necessarily be identical;
@@
    in the general case (by accounting for arbitrary strings beyond a handful of pre-selected examples), the expected compression rate (or expected number of factors in their respective parsings) would not be correlated at the asymptotic limit (as the length of strings/objects increases), even though they may not actually be for an a priori chosen set of particular examples (especially for short sequences).
@@
For the Molecular Assembly Index (MA) introduced in [1] as an application of Ai within the context of molecular complexity, we show that it is not only highly correlated with the LZW of InChI strings, but also highly correlated with the length of InChI strings themselves. In addition, we showed that one of the main benefits the authors claim about Ai (‘molecular assembly’ or MA) as being experimentally measurable using mass spectrometry data, is also observed when applying LZW directly on the InChI chemical compound descriptions, implying that it is neither something of particular interest to report nor something only Ai can do.
In order to expose (a) as false, it is always possible to pick particular cases of strings for which the parsing, according to LZ77, differs from the one according to LZ78 [25]. These are 2 different LZ compression schemes that differ trivially yet they are based on the same principles, and both are derived and belong to the same family.
@@
One can see in [4, Figure 3] that the authors omit the dictionary built for the right square (i.e., the one for the parsing according to the assembly index calculation method). Specifically, this assembly-coding dictionary is
$(1,\textnormal{``z''}),$ (7)
$(2,[1,1]),$
$(3,[2,2]),$
$(4,[3,3])$
@@
The final object would be given by the codeword list $[4,4]$ as indicated by the right square in [4, Figure 3]. Notice that the set of vertices of the minimum rooted assembly subspace of the final object is given by the set $\{1,2,3,4,[4,4]\}$ composed of pointers; and the assembly index is the size of this dictionary excluding the basis element under pointer $1$, that is, Ai =51=4=5-1=4. Another fact overlooked by the authors is that if one uses an LZ factorisation/parsing scheme straightforwardly defined by adding to the dictionary the longest prefix of the subsequent subsequence that has already occurred, or a single symbol (in our case, “z”) if there is no such nonempty prefix, then one would obtain exactly the desired parsing
$\textnormal{z}|\textnormal{z}|\textnormal{zz}|\textnormal{zzzz}|\textnormal{zzzzzzzz}$ (8)
@@
that the authors of AT overlook, a parsing which gives the exact assembling process in the minimum rooted assembly subspace of the final object in Equation (8). Notice that the number of factors/codewords in Equation (8) (minus 1) is the assembly index of the final object. Except for the initial basis element “z”, the parsing in Equation (8) exactly corresponds to the sequence of edge labels of the rooted assembly path of the final object. In this trivial example, we see that the shortest rooted assembly path and the longest one (which belong to the minimum rooted assembly subspace) are the same.
This also answers the AT authors’ challenge and questioning in [4] regarding the lack of dictionary or memory in Ai, this is yet another fallacy, the dictionary is in the parsed molecules themselves. It is not that they do not. and that LZ algorithms are artificial computational constructs, it is just that the authors have not been capable of finding the obvious technical analogies and equivalences that in computer science are separated for implementation or pedagogical purposes.
@@
Regarding (b), the authors’ counterargument constitutes a combination of the “Texas sharp shooter” and the “hasty generalisation” fallacies: where one takes the global properties as always the same as the properties resulting from analyses of particular local cases, which they would invariably be, given that the criteria for the latter analyses were chosen to produce the desired local properties.
@@
From the equivalence between Ai and dictionary-based compression schemes formally demonstrated in [17], we know that the set of vertices/objects of the minimum subspace encodes/compresses the final object (in our case of interest, the input string). This holds whether the basis elements are included or not, since the compression rate obtained in the former case or the latter case differs only by an independent constant. Thus, the number of codewords in the list (which results from the parsing of the string according to the assembly index method) is equivalent to the length of the longest (rooted) assembling paths in the minimum subspace, except for those that are basis elements. In other words, they show all the elements of a typical statistical compression algorithm as proved to be.
@@
For example, for the input string “abracadabra”, the dictionary derived overlooked by the authors of AT in their example calculations is:
‘b’ $:\textnormal{`!'},$ (9)
‘r’ $:\textnormal{`@'},$
‘a’ $:\textnormal{`\textdollar'}$,
‘c’ $:\textnormal{`\%'},$
‘d’ $:\textnormal{`\&'},$
‘ab’ $:\textnormal{`('},$
‘abr’ $:\textnormal{`)'},$
‘abra’ $:\textnormal{`*'}$
@@
Thus, Ai is simply 81=78-1=7, which is given by the number of single symbols inside the operator $\cdot$] that represent an edge in a rooted assembling path whose label is taken from one of the vertices of the minimum subspace.
@@
‘b’ $:\textnormal{`000'},$ (11)
‘r’ $:\textnormal{`001'},$
‘a’ $:\textnormal{`010'},$
‘c’ $:\textnormal{`011'},$
‘d’ $:\textnormal{`100'},$
‘ab’ $:\textnormal{`101'},$
‘br’ $:\textnormal{`110'},$
‘ra’ $:\textnormal{`111'},$
‘ac’ $:\textnormal{`0000'},$
‘ca’ $:\textnormal{`0001'},$
‘ad’ $:\textnormal{`0010'},$
‘da’ $:\textnormal{`0011'},$
‘abr’ $:\textnormal{`0100'}$
while the codeword list is
@@
In the case of LZW/LZ78 we know that the size in bits of the compressed form of the input string (in the asymptotic limit) grows under $c_{1}(n)\mathbf{O}(\log(c_{1}(n)))$ [25, 35], where $c_{1}(n)$ is the size of the LZW/LZ78 dictionary for a string of length $n$. The same also holds for the compressed form (with size $K_{Ai}(x)$ in bits) of the input string $x$ according to the assembly index calculation method, so that
$K_{Ai}(x)\leq c_{2}(n)\mathbf{O}(\log(c_{3}(n)))\leq c_{2}(n)\mathbf{O}(\log(c_{2}(n)))\text{ ,}$ (13)
where $c_{2}(n)$ is the length of the list of codewords that constitutes the string $x$ and $c_{3}(n)=\mathrm{Ai}(n)$ is the assembly index that is the size of the Ai dictionary [17] for a string $x$ of length $n=|x|$ with $\mathrm{Ai}(n)\leq c_{2}(n)$. This shows that Ai does implement an implicit dictionary in accordance with the expectation of belonging to the family of popular statistical compression algorithms LZ that is widely used in complexity science.
Sequences with larger values of $c_{i}(\cdot)$, i.e., the number of distinct phrases or codewords in the dictionary, to which the pointers necessarily recur less often, would have a lower probability. This is the same key idea used to demonstrate the optimality of LZ schemes toward the limit given by the noiseless coding theorem; and it is the same idea behind the assembly index as a complexity measure, which the authors claim is a novelty.
When it comes to molecules, the dictionaries are simply embedded in the molecules themselves. Every time their representations are parsed and counted, a dictionary is actually being built behind the operation of the Ai metric.
Here we further investigate the relationships between these values with experiments that generalise the purported counterexamples to much longer sequences of different configurations, as explained in Section 4.2, thus demonstrating that (b) is indeed such a fallacy.
4.1 Stochasticity and Connections of BDM with Shannon Entropy and LZW
In its expanded form, the BDM of an object $x$, according to the partition $i$, is defined by
$BDM(x,i,m)=\sum_{(r_{j},n_{j})\in P_{i}(x)}(\log(n_{j})+K_{m}(r_{j}))\text{ ,}$ (14)
@@ 4.1 Stochasticity and Connections of BDM with Shannon Entropy and LZW
  • $P_{i}(x)$ is the set of pairs $(r_{j},n_{j})$ obtained when decomposing the object $x$ according to a partition $i$ of contiguous parts $r_{j}$, each of which appears $n_{j}$ times in such a partition (i.e., $n_{j}$ is the multiplicity of $r_{j}$).
@@
Notice that Equation (14) can be expressed for many particular cases, e.g. in the case of unidimensional objects (such as bitstrings) but also any other object where $m$ indicates the method given by the pre-computed table of complexity estimations derived from CTM [6, 28, 36, 31].
Also notice that if one takes a partition $i_{0}$ for which the object’s size is the partition size, then $BDM(x,i_{0},m)$ becomes equivalent to the approximation method indexed as $m$, e.g. the CTM. That is, there is $m$ such that
@@
where $Q$ is a computable (proper) subset of $\{\mathbf{M}:\exists w^{\prime}(\mathbf{M}(\emptyset)=w^{\prime})\}$ and $Q(x)\subset\{\mathbf{M}:\mathbf{M}(\emptyset)=x\}$ a computable subset of $Q$ that the method $m$ calculates, where the algorithmic probability (AP) of $x$ is approximated by
@@
and $\mathbf{M}(x)=y$ denotes algorithm $\mathbf{M}$ returning $y$ with input $x$.
Deriving from this equivalence, one of the properties of BDM is that it is a complexity measure that approaches $\mathbf{K}(\cdot)$ in the best case scenario, and approaches block (Shannon) Entropy in the worst case [6], that is,
[…]
The left inequalities of Equation (16) always hold, from the basic properties of algorithmic complexity. From classical information theory, the rightmost inequality holds for sufficiently large $|x|$ and any computable compression method $m^{\prime}$ that is capable of matching (or surpassing) the statistical compression/encoding achieved by, e.g., Huffman or LZW for each codeword, where
$\mathbf{H}_{i}(X^{i})=-\sum_{(r_{j},n_{j})\in P_{i}(x)}p(r_{j})\log(p(r_{j}))$ (17)
is the block (Shannon) Entropy $\mathbf{H}_{i}$ of an i.i.d. source $X^{i}$ such that $p(r_{j})\to\frac{n_{j}}{N_{i}}$ as $|x|\to\infty$, the random variable $X^{i}$ can assume values in the set $\{r_{1},\dots,r_{j},\dots,r_{|P_{i}(x)|}\}$, and $N_{i}=\sum_{(r_{j},n_{j})\in P_{i}(x)}n_{j}$.
Thus, in case $m^{\prime}$ is a sufficiently powerful computable approximation and $i$ a partition closer to the partition $i_{0}$, one obtains $\mathbf{K}(x)\sim BDM(x,i,m^{\prime})$; otherwise, in case the object $x$ is indeed generated by a stochastic source, one obtains $BDM(x,i,m^{\prime})\leq\mathbf{O}(1)+N_{i}\mathbf{H}_{i}(X^{i})$ approaching the equality the more statistically random the source is. Also note that, as demonstrated in [6] for regular bidimensional partitions, an algorithm can always losslessly retrieve $x$ from the information in the parts/blocks calculated in $BDM(x,i,m^{\prime})$ by an encoding scheme $\left<\cdot\right>_{BDM}$ such that
$|\left<x\right>_{BDM}|\leq\mathbf{O}(1)+BDM(x,i,m^{\prime})+\mathbf{O}(N_{i}\log(|P_{i}(x)|))$ (18)
@@
This means that for objects generated according to a probability distribution sufficiently close to the uniform distribution and a sufficiently powerful $m^{\prime}$, BDM can always achieve a statistical lossless compression of the same order as $\mathbf{H}$, while the value of $BDM(x,i,m^{\prime})$ can be smaller than $N_{i}\mathbf{H}_{i}(X^{i})$. Therefore, $\mathbf{H}$ is a (statistical) compression method subsumed within BDM.
A false argument repeated by the authors of Assembly Theory is that approximations to algorithmic Kolmogorov complexity are uncomputable, unlike assembly theory [26]. Not only is this false, as BDM is fully computable [37, 38], but also, the assembly index itself is an approximation to $\mathbf{K}$ equivalent to $\mathbf{H}$ and LZ compression, as we have proven [17], therefore a loose upper bound on $\mathbf{K}$ itself. Hence, the assembly index offers no advantage.
BDM can also deal with stochasticity, by exploring all possible computer program pathways that assemble an object, as fully explained in Section 3.5.